12 year old ben's grandfather is 66. How many years ago was his grandfather 10 times older than ben?
Answer by robertb(5830)

10(12 - x) = 66- x,
120 - 10x = 66-x,
54 = -x + 10x,
54 = 9x,
x = 6. Therefore 6 years ago Ben's grandfather was 10 times as old as Ben.